The perimeter of a rectangle is 42cm and its area is 68cmsq. Find its breadth and lenght in square meter?
Answer by ewatrrr(24785)

P = 2L + 2w = 42cm
L + w = 21cm
(21-w)w = 68cm^2
w^2 - 21w + 68 = 0
(w - 17)(w-4) = 0 , w = 4cm,17cm
Rectangle is 4cm by 17cm